# Entanglement Microscopy: Tomography and Entanglement Measures via Quantum Monte Carlo

### v3

@@ Results
We can actually obtain the scaling analytically in the large $T$ limit for general models and separations. In order to determine $\mathcal{E}^{f}$ in a thermal state $\rho=\exp(-\beta H)/\mathcal{Z}$, we need to find how the fermion Green’s function $g_{ij}=\operatorname{Tr}(c^{\dagger}_{i}c_{j}\rho)$ decays to zero. We will then have $\mathcal{E}^{f}=\alpha|g_{ij}|^{2}+\cdots$, where the coefficient $\alpha$ depends on the filling, and the ellipsis denote subleading terms. By expanding $\exp(-\beta H)=\sum_{n}(\beta H)^{n}/n!$ we see that the first non-zero contribution to the Green’s function will come from $\operatorname{Tr}(H^{n_{r}}c^{\dagger}_{i}c_{j})$ such that the hopping term of $H^{n_{r}}$ connects sites $i$ and $j$. The Green’s function thus decays as $1/T^{n_{r}}$, and we have
@@
Finally, Fig. 3 (b) shows the scaling of $\mathcal{E}^{f}$ as we increase the separation between the two sites at the GNY QCP. In contrast to $\mathcal{E}^{b}$ for the Ising QCP, the fermionic LN shows long-range entanglement, manifested by its power-law decay with $r$. The dashed line shows the CFT prediction [9] with power $4\Delta_{\psi}\approx 4.174$, where the fermion scaling dimension was obtained via conformal bootstrap [43]. The algebraic tail of the fermionic LN follows from the general structure of the space of separable states in fermion systems owing to the parity superselection rule [10].
@@
Figure 4: Searching for genuine tripartite entanglement in the 1d/2d quantum Ising models. In 1d, the criterion $W_{1}\geq 0$ detects GTE for 3 consecutive spins at all the values of $h$. In 2d for 3 spins in a square plaquette, $W_{1}\leq 0$, both in the small size exact diagonalization (ED) and large-scale QMC results, signifying the absence of the GTE in a large region near the QCP. Similar behavior for $W_{2}$ is shown in the SM [36]. All QMC calculations are done with $\beta=L$.
@@
Entanglement witnesses. In 1d, GTE of three consecutive spins was found in the 1d anisotropic XY model, which includes the Ising case as a special limit [13], using a quantum $k$-separability criterion [60]. GTE has also been studied in a cluster-Ising model that gives rises to RDMs with special fine-tuned properties [61], while multi-partite entanglement of the entire system, not subregions as considered here, was studied for small system sizes in the TFIM in dimensions 1 to 3 [62]. Quantum microscopy of the GTE beyond 1d has not been previously performed. Fig. 4 shows the GTE criterion $W_{1}$ for the three closest sites as a function of the transverse field $h$ for both the 1d and 2d TFIMs. The 1d data, obtained from the exact RDM via a Jordan-Wigner transformation [63], shows that GTE exists for all finite values $h$, and becomes maximal near the QCP. In fact, $dW_{1}/dh$ diverges at QCP, exactly as was found for a related criterion $dI_{2}/dh$ in [13]. In sharp contrast, in the 2d case, the large-scale QMC data has $W_{1}\leq 0$ for all values $h$, indicating that no GTE is detected. We have also evaluated another complementary criterion $W_{2}$, and reached the same conclusion in the range $20<h/h_{c}<$ 2 (see SM [36]), although more powerful criteria are needed to make a definitive conclusion. Furthermore, we have checked that in both 1d and 2d no GTE is detected if the sites are non-adjacent. Extrapolating to higher dimensions, we expect that GTE decreases with increasing $d$ partly due to the monogamy of entanglement. One could test this in the 3d TFIM on the cubic lattice.
@@ Discussion
In this work, we have put forward a framework – entanglement microscopy – to perform full tomography within QMC, giving access to the true entanglement hidden in quantum many-body systems. To demonstrate the power of our scheme, we study the phase diagram of two representative QCPs in 2 spatial dimensions: the transverse field Ising model and the GNY chiral-Ising transition of Dirac fermions. We find i) the Ising QCP exhibits short-range entanglement with a finite sudden death as a function of separation and temperature; ii) the fermionic QCP exhibits power-law decay of the fermionic LN at large separations between the subregions; iii) there is no detectable 3-party entanglement with our witnesses in a large window near the 2d Ising QCP, while it is present in the 1d counterpart. Our results demonstrate that controlled entanglement microscopy is now possible on large-scale quantum many-body systems in any dimension. A new window into quantum many-body entanglement has opened, with countless systems waiting to be explored.
In the future, it will be interesting to explore the multipartite entanglement at various exotic QCPs, including transitions into a non-Fermi-liquid [64, 65], deconfined QCPs for both fermion and spin models [66, 67, 68, 69, 70, 71, 72]. In such explorations, the recent realisation that the very structure of the space of physical states determines the fate of entanglement as a system evolves will be useful to guide entanglement microscopy in realistic many-body Hamiltonians [10].
@@ Methods
In the usual SSE simulation [30], the path-integral is Taylor expanded as
$\begin{split}Z&=\operatorname{Tr}\{e^{-\beta H}\}=\sum_{\psi}\bra{\psi}e^{-\beta H}\ket{\psi}\\ &=\sum_{\{b_{1},b_{2}\cdots b_{M}\}}\frac{\beta^{n}(M-n)!}{M!}\sum_{\psi}\bra{\psi}H_{b_{M}}\cdots H_{b_{2}}H_{b_{1}}\ket{\psi},\end{split}$ (7)
where $\{b_{1},b_{2}\cdots b_{M}\}$ denotes the type of operators on $M$ imaginary time slices. The sampling of bonds and states is carried out by performing diagonal update, which add and remove diagonal operators, and loop update, which interchange diagonal and off-diagonal operators.
In sampling RDM in SSE, where imaginary time boundary condition is opened for sites in region $A$, one simply break the loop at these points in loop update, and the detailed balance condition will be satisfied. Hence, the frequency of any edge configuration is proportional to the corresponding element in the RDM.
Determinant quantum Monte Carlo. For the $t-V$ model, we employ the determinant quantum Monte Carlo (DQMC) scheme [31, 32, 33], where the fermion interaction is decoupled into fermion bilinear coupled to auxiliary field. In each configuration, the free fermionic Green’s function can be calculated, and higher order correlation is evaluated using Wick’s theorem.
@@ Acknowledgments
"Entanglement Microscopy: Tomography and Entanglement Measures via Quantum Monte Carlo"
@@
In finding the QCP of $t-V$ model, we perform the finite-temperature DQMC in various system sizes and set $\beta$ scale linearly with system size, since the fermion dispersion is linear ($z=1$) at the QCP. Fig. S5 shows the order parameter $M^{2}=S(0,0)$, where $S(k_{x},k_{y})=\sum_{i,j,\alpha}e^{-i(k_{x},k_{y})\cdot{\mathbf{r}_{ij}}}(n_{i,\alpha}n_{j,\alpha}-n_{i,\alpha}n_{j,\bar{\alpha}})$, correlation ratio $R=1-\frac{S(2\pi/L,0)}{S(0,0)}$, the finite-size crossing of them and data collapse of $M^{2}$ using the critical exponents from Ref. [43].

### v4

@@ Results
As temperature increases, the nearest neighbor $\mathcal{E}^{f}$ does not suffer a sudden death. Instead, it decays algebraically as shown in Fig. 2 (b) and Fig. 3 (d). We can actually obtain the scaling analytically in the large $T$ limit for general models and separations. In order to determine $\mathcal{E}^{f}$ in a thermal state $\rho=\exp(-\beta H)/\mathcal{Z}$, we need to find how the fermion Green’s function $g_{ij}=\operatorname{Tr}(c^{\dagger}_{i}c_{j}\rho)$ decays to zero. We will then have $\mathcal{E}^{f}=\alpha|g_{ij}|^{2}+\cdots$, where the coefficient $\alpha$ depends on the filling, and the ellipsis denote subleading terms. By expanding $\exp(-\beta H)=\sum_{n}(\beta H)^{n}/n!$ we see that the first non-zero contribution to the Green’s function will come from $\operatorname{Tr}(H^{n_{r}}c^{\dagger}_{i}c_{j})$ such that the hopping term of $H^{n_{r}}$ connects sites $i$ and $j$. The Green’s function thus decays as $1/T^{n_{r}}$, and we have
@@
Finally, Fig. 3 (b) shows the scaling of $\mathcal{E}^{f}$ as we increase the separation between the two sites at the GNY QCP. In contrast to $\mathcal{E}^{b}$ for the Ising QCP, the fermionic LN shows long-range entanglement, manifested by its power-law decay with $r$. The dashed line shows the CFT prediction [21] with power $4\Delta_{\psi}\approx 4.174$, where the fermion scaling dimension was obtained via conformal bootstrap [46]. The algebraic tail of the fermionic LN follows from the general structure of the space of separable states in fermion systems owing to the fermion parity superselection rule [22].
Genuine Tripartite Entanglement
@@
Figure 4: Searching for genuine tripartite entanglement in the 1d/2d quantum Ising models. (a) In 1d, the criterion $W_{1}\geq 0$ detects GTE for 3 consecutive spins at all the values of $h$. In 2d for 3 spins in a square plaquette, $W_{1}\leq 0$, both in the small size exact diagonalization (ED) and large-scale QMC results, signifying no detection of GTE in a large region near the QCP. Similar behavior for $W_{2}$ is shown in the SM [38]. All QMC calculations are done with $\beta=L$. Panel (b) shows the logarithmic divergence of the derivative $\frac{dW_{1}}{dh}$ near QCP for the 1d $L\!=\!\infty$ data: $\frac{dW_{1}}{dh}|_{h}=\frac{W_{1}(h+\Delta h)-W_{1}(h-\Delta h)}{2\Delta h}$, with step size $\Delta h=10^{-4}$.
@@
In 1d, GTE of three consecutive spins was studied in the 1d anisotropic XY model, which includes the Ising case as a special limit [27], using a quantum $k$-separability criterion [59]. GTE has also been studied in a cluster-Ising model that gives rise to RDMs with special fine-tuned properties [60], while multi-partite entanglement of the entire system, not subregions as considered here, was studied for small system sizes in the TFIM in dimensions 1 to 3 [61]. GTE beyond 1d has not been previously computed. Fig. 4 shows the GTE criterion $W_{1}$ for the three closest sites as a function of the transverse field $h$ for both the 1d and 2d TFIMs. The 1d data, obtained from the exact RDM via a Jordan-Wigner transformation [62], shows that GTE exists for all finite values $h$, and becomes maximal in the paramagnetic phase near the QCP. The absence of sudden death with $h$, can be explained similarly to the LN case above. The asymptotic state at $h=\infty$ is a pure product state, and so is the RDM of 3 spins. This state actually lies at the boundary of the all biseparable states since we can perturb with an arbitrarily small $\epsilon$: $\sqrt{1-\epsilon}|\uparrow\uparrow\uparrow\rangle+\sqrt{\epsilon}|\downarrow\downarrow\downarrow\rangle$ to generate a cat / GHZ state (with unequal amplitudes), which possesses GTE as detected by $W_{1}$. At small $h$, we have a rank 2 RDM (since the full GS is the ferromagnetic cat state), which is rank deficient. In the zero-rank subspace, we can perturb with a $W$-state: $(1-\epsilon)\rho+\epsilon|W\rangle\langle W|$ where $|W\rangle=(|\downarrow\uparrow\uparrow\rangle+|\uparrow\downarrow\uparrow\rangle+|\uparrow\uparrow\downarrow\rangle)/\sqrt{3}$. The $W_{2}$ criterion detects GTE when $\epsilon>0$. This shows that the asymptotic state is on the boundary of biseparable states, and so the entanglement is only lost at $h=\infty$.
In sharp contrast, in the 2d case, the large-scale QMC data has $W_{1}\leq 0$ for all values $h$, indicating that no GTE is detected. We have also evaluated another complementary criterion $W_{2}$, and reached the same conclusion in the range $20<h/h_{c}<$ 2 (see SM [38]), although more powerful criteria are needed to make a definitive conclusion. Furthermore, we have checked that in both 1d and 2d no GTE is detected if the sites are non-adjacent. Extrapolating to higher dimensions, we expect that GTE decreases with increasing $d$ partly due to the monogamy of entanglement. One could test this in the 3d TFIM on the cubic lattice.
@@
Scaling of multipartite entanglement near criticality
Let us now look more closely at the scaling of entanglement measures near the transition. We shall first present an analysis that establishes the singular scaling of $W_{1}$ in a large class of quantum critical points that includes the Ising CFT in $d\geq 1$, and then give a general argument for other multipartite entanglement measures (which includes bipartite as a special case). Let $h$ be the non-thermal parameter that tunes the system to a QCP. Let $U_{c}$ be the optimal LU transformation that yields $W_{1}$ at $h_{c}$, and $\check{\rho}=U_{c}\rho U_{c}^{\dagger}$ the transformed RDM. As $h$ is varied away from $h_{c}$, by the Envelope Theorem [63], the only contribution comes from the direct variation of the RDM, with $U_{c}$ unchanged, see Supplemental Material [38] for a derivation. The density matrix elements are linear combinations of $n$-point functions of Pauli operators with $n\leq 3$. The Pauli strings with non-zero expectation value need to preserve the symmetries of the finite-size Hamiltonian, which we take to have a unique groundstate. These operators will generically have an overlap with the relevant low energy scalar of the quantum critical theory, which we call $\varepsilon$, with (spatial) scaling dimension $\Delta_{\varepsilon}=d+z-1/\nu$, where $z$ is the dynamical critical exponent, and $\nu$ is the correlation length exponent. As such, the dominant variation of the expectation value at small $|h-h_{c}|$ will be proportional to $|h-h_{c}|^{\Delta_{\varepsilon}\nu}$. In turn, this will lead to a singular variation $W_{1}=W_{1}(h_{c})+a|h-h_{c}|^{\Delta_{\varepsilon}\nu}+\cdots$.
A more basic argument can be made to explain the singularity of a generic entanglement measure or criterion, which we call $\mathcal{M}$, irrespective of the number of spins / parties considered. $\mathcal{M}$ depends on the expectation values within the subregion, and in a symmetric groundstate the relevant scalar will typically dominate those, yielding the singular derivative:
$\frac{d\mathcal{M}}{dh}\Big{|}_{h\to h_{c}}=\alpha\,|h-h_{c}|^{\Delta_{\varepsilon}\nu-1}$ (7)
It may happen that $\alpha$ vanishes so that the singular behavior appears in higher-order derivatives, as is the case for the next-nearest neighbor concurrence in the 1d TFIM [25], but that situation is not generic.
In the case of the 1d TFIM, described by the 1d Ising CFT, we have $\Delta_{\varepsilon}=\nu=z=1$, and a logarithmic divergence appears instead:
$\frac{dW_{1}}{dh}\Big{|}_{h\to h_{c}}=-\alpha\ln|h-h_{c}|$ (8)
where $\alpha\approx 0.097$, as can be seen in Fig.4b. The GS correlation functions have indeed a logarithmic $\ln|h-h_{c}|$, as can be verified from the exact solution [64]. The general result (7) also explains the logarithmic divergence numerically seen in various genuine 3-spin and 4-spin entanglement metrics in the 1d TFIM [27, 65]. In the 2d Ising CFT, $\Delta_{\varepsilon}\nu=0.890$, such that the exponent in (7) is $-0.11$. Such a divergence is difficult to reveal using our finite-size data, but the derivative becomes maximal in close proximity to the QCP, see Fig. S5 in the Supplemental Materials [38].
It is worth noting that in cases of symmetry breaking, which can occur for instance in the magnetic phase of the quantum Ising model in the thermodynamic limit, the critical exponent in the ordered phase will generically differ from that in the symmetric phase. In the symmetry broken phase, $\varepsilon$ should then be replaced by the lowest operator that acquires an expectation value. In the Ising case, this will be the order parameter $\phi$, so (7) will depend on $\Delta_{\phi}$ instead of $\Delta_{\varepsilon}$ when $h<h_{c}$.
@@ Discussion
In this work, we performed entanglement microscopy to access the true entanglement hidden in quantum many-body systems, by scrutinizing the full RDMs of microscopic subregions obtained by ED and QMC [24]. We have studied the phase diagram of two representative QCPs in 1 and 2 spatial dimensions: the transverse field Ising model, both in 1d and 2d, and the 2d GNY chiral-Ising transition of Dirac fermions. We found i) the Ising QCP exhibits short-range entanglement with a finite sudden death as a function of separation and temperature; ii) the fermionic QCP exhibits power-law decay of the fermionic LN at large separations between the subregions; iii) there is no detectable 3-party entanglement with our witnesses in a large window near the 2d Ising QCP, while it is present in the 1d counterpart. Finally, we present general non-pertubative results regarding the scaling of multipartite entanglement measures near quantum critical points. Our results demonstrate that precise entanglement microscopy is possible on large-scale quantum many-body systems beyond 1d. A new window into quantum many-body entanglement is opening, with countless systems waiting to be explored.
In the future, it will be interesting to explore the multipartite entanglement at various exotic QCPs, including transitions into a non-Fermi-liquid [66, 67], deconfined QCPs for both fermion and spin models [9, 68, 69, 70, 10, 71, 72], and symmetric mass generation transitions [73].
@@ Methods
In the SSE simulation [74], the partition function $Z$ is Taylor-expanded as a power series of Hamiltonian operators, and each term is sampled stochastically. As mentioned in the previous section, the RDM matrix element resembles a partition function but with imaginary time boundary conditions open in the region $A$, which is expressed as
$\begin{split}\langle\alpha|\rho_{A}|\tilde{\alpha}\rangle&=\frac{1}{Z}\sum_{\xi}\bra{\alpha,\xi}e^{-\beta H}\ket{\tilde{\alpha},\xi}\\ &=\frac{1}{Z}\sum_{\xi}\sum_{n}\frac{\beta^{n}}{n!}\bra{\alpha,\xi}H^{n}\ket{\tilde{\alpha},\xi}\end{split}$ (9)
after series expansion of $e^{-\beta H}$. Therefore, the sampling is almost identical to the standard SSE, but now we simply break the imaginary-time periodicity for loops that pass through sites in the region $A$ when doing loop updates. The frequency of any configuration pair $(\alpha,\tilde{\alpha})$ is proportional to the magnitude of the corresponding element in the RDM, i.e., $\bra{\alpha}\rho_{A}\ket{\tilde{\alpha}}$, according to the detailed balance condition. That is,
$\frac{\bra{\alpha}\rho_{A}\ket{\tilde{\alpha}}}{\bra{\alpha^{\prime}}\rho_{A}\ket{\tilde{\alpha}^{\prime}}}=\frac{P(\alpha,\tilde{\alpha})}{P(\alpha^{\prime},\tilde{\alpha}^{\prime})},$ (10)
with $P$ denotes the probability of pair of states $(\alpha,\tilde{\alpha})$ to appear. Interested readers are welcome to read Ref. [24] for more details.
Determinant quantum Monte Carlo. For the $t-V$ model, we employ the determinant quantum Monte Carlo (DQMC) scheme [29, 30, 31], where the fermion interaction is decoupled into fermion bilinear coupled to the auxiliary field. In each configuration, the free fermionic Green’s function can be calculated, and higher-order correlation functions are evaluated using Wick’s theorem.
Data availability
The data that support the findings of this study are available from the corresponding author upon reasonable request.
Code availability
All numerical codes in this paper are available upon reasonable request to the authors.
@@ Acknowledgments
Author contributions
T.-T.W. and M.H.S. realized the RDM sampling algorithm. T.-T.W. performed the QMC simulation for fermion $t-V$ model. M.H.S. performed the QMC simulation for TFIM. L.L. calculated the exact result in 1d TFIM and performed the ED calculation in 2d TFIM. All the authors contributed to the analysis of the results. W.W.-K. and Z.Y.M. supervised the project.
Additional information
Supplementary Information
is available in the online version of the paper.
Competing interests: The authors declare no competing interests.
@@
"Entanglement Microscopy: Tomography and Entanglement in Many-Body Systems"
@@
Figure S5: Derivative of $W_{1}$ with respect to h for the ED 5x5 Square Lattice.
Building upon our discussion in the main text, we now present the derivative of $W_{1}$ with respect to $h$ for the 2D system with $L=5$ using ED, as shown in Fig. S5. The derivative exhibits a pronounced peak in close proximity to the quantum critical point (QCP), reinforcing the main text’s observation that the derivative becomes maximal near the QCP.
@@
Derivation of $dW_{1}/dh$
In this section, we aim to give a general formula for the derivative of any genuine multipartite entanglement (GME) measure $W$, which includes both $W_{1}$ and $W_{2}$, and depends on some optimization procedure. This problem is part of a broader class of problems involving the differentiation of a maximum value function, which is addressed by the well-known Envelop Theorem [63]. Specifically, we will demonstrate that the derivative $\frac{dW_{1}}{dh}$ exhibits a linear dependence on $\frac{d\rho}{dh}$, achieved by analyzing the behavior of the objective function $W(h)$ under variations in the parameter $h$.
Consider $\rho(h)$ parametrized by a single parameter $h$. The local unitary transformation $U=U_{1}\otimes U_{2}\otimes U_{3}$ is parametrized by a set of angles $x=\{x_{i}\}$. The objective function is defined as $f(x;h)$. Using this, we define the quantity $W(h)$ as:
$W(h)=\max_{x}f(x;h)$
The optimal set of angles $x^{*}$ must satisfy the condition:
$\frac{\partial f}{\partial x_{i}}(x^{*})=0\quad\forall i$
Using these optimal angles $x^{*}(h)$, the function $W(h)$ can be rewritten as:
$W(h)=f(x^{*}(h),h)$
Now, taking the derivative of $W(h)$ with respect to $h$, we obtain:
$\frac{dW}{dh}=\sum_{i}\frac{\partial f(x^{*})}{\partial x_{i}}\frac{\partial x_{i}^{*}}{\partial h}+\frac{\partial f}{\partial h}$
Since the optimal angles satisfy $\frac{\partial f}{\partial x_{i}}(x^{*})=0$, the first term vanishes, leaving us with:
$\frac{dW(h)}{dh}=\frac{\partial f(x^{*}(h),h)}{\partial h}$
Next, we apply these results to the specific case of $W_{1}$ by defining the function $g(\rho)$ as:
$g(\rho)=\left|\rho_{1,8}\right|-\sqrt{\rho_{2,2}\rho_{7,7}}-\sqrt{\rho_{3,3}\rho_{6,6}}-\sqrt{\rho_{4,4}\rho_{5,5}}$
The density matrix $\tilde{\rho}$ after the unitary transformation $U(x)$ is given by:
$\tilde{\rho}=U(x)\rho(h)U^{\dagger}(x)$
Thus, the objective function for $W_{1}$ becomes:
$f(x;h)=g(\tilde{\rho})$
We now take the derivative of $W_{1}(h)$ with respect to $h$:
$\frac{dW_{1}}{dh}=\frac{\partial g}{\partial h}=\sum_{ij}\frac{\partial g}{\partial\tilde{\rho}_{ij}}\frac{\partial\tilde{\rho}_{ij}}{\partial h}$
Since $\tilde{\rho}=U(x^{*})\rho(h)U^{\dagger}(x^{*})$, we can apply the chain rule to write:
$\frac{\partial\tilde{\rho}}{\partial h}=U(x^{*})\frac{\partial\rho(h)}{\partial h}U^{\dagger}(x^{*})$
Substituting this into the expression for $\frac{dW_{1}}{dh}$, we get:
$\frac{dW_{1}}{dh}=\sum_{ij}\frac{\partial g}{\partial\tilde{\rho}_{ij}}\left[U(x^{*})\frac{\partial\rho(h)}{\partial h}U^{\dagger}(x^{*})\right]_{ij}$
Finally, introducing the optimal unitary transformation $U_{c}=U(x^{*})$, the result simplifies to:
$\frac{dW_{1}}{dh}=\sum_{ij}\frac{\partial g}{\partial\tilde{\rho}_{ij}}\left[U_{c}\frac{\partial\rho(h)}{\partial h}U_{c}^{\dagger}\right]_{ij}$
At this point, we have arrived at the key result:
$\frac{dW_{1}}{dh}=\sum_{ij}\frac{\partial g}{\partial\rho_{ij}}\left[U_{c}\frac{\partial\rho(h)}{\partial h}U_{c}^{\dagger}\right]_{ij}$
where $U_{c}$ is the optimal unitary matrix. We have thus shown that the derivative $\frac{dW_{1}}{dh}$ depends linearly on $\frac{d\rho}{dh}$, as required.
@@
In finding the QCP of $t-V$ model, we perform the finite-temperature DQMC in various system sizes, setting $\beta$ to scale linearly with system size, since the fermion dispersion is linear ($z=1$) at the QCP. Fig. S6 shows the order parameter $M^{2}=S(0,0)$, where $S(k_{x},k_{y})=\sum_{i,j,\alpha}e^{-i(k_{x},k_{y})\cdot{\mathbf{r}_{ij}}}(n_{i,\alpha}n_{j,\alpha}-n_{i,\alpha}n_{j,\bar{\alpha}})$, correlation ratio $R=1-\frac{S(2\pi/L,0)}{S(0,0)}$, the finite-size crossing of them and data collapse of $M^{2}$ using the critical exponents from Ref. [46].